# Fourier Learning Machines: Nonharmonic Fourier-Based Neural Networks for Scientific Machine Learning

### v1

@@ 3 Mathematical Preliminaries
$\widehat{f}_{H}(\boldsymbol{x})=\sum_{\boldsymbol{n}\in\mathbb{Z}^{m}}c_{\boldsymbol{n}}\,e^{i\boldsymbol{n}\cdot\boldsymbol{x}},$ (5)
where $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{m})\in[-\pi,\pi]^{m}$ , $\boldsymbol{n}=(n_{1},n_{2},\dots,n_{m})\in\mathbb{Z}^{m}$ and $\boldsymbol{n}\cdot\boldsymbol{x}$ represents their dot product. The analytical expression for calculating the Fourier coefficients is
$c_{\boldsymbol{n}}=\frac{1}{(2\pi)^{m}}\int_{[-\pi,\pi]^{m}}f(\boldsymbol{x})\,e^{-i\boldsymbol{n}\cdot\boldsymbol{x}}\,d\boldsymbol{x},$
@@
$\widehat{f}_{H}(\boldsymbol{x})=\sum_{\boldsymbol{n}\in\mathbb{N}_{0}^{m}}\Bigl{[}$ $a_{1}^{(\boldsymbol{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\cos(n_{m}x_{m})$ (6)
$+$ $a_{2}^{(\boldsymbol{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\sin(n_{m}x_{m})+\ \dots$
$+$ $a_{2^{m}}^{(\boldsymbol{n})}\,\sin(n_{1}x_{1})\sin(n_{2}x_{2})\dots\sin(n_{m}x_{m})\Bigr{]},$
@@
for input $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{m})\in[-\pi,\pi]^{m}$ and frequency components $\boldsymbol{n}=(n_{1},n_{2},\dots,n_{m})\in\mathbb{N}_{0}=\{0,1,2,\dots\}$. This trigonometric separable form has the advantage that the set of basis functions, associated with $\boldsymbol{n}$, simultaneously accounts for all possible $2^{m}$ symmetric frequency components. Here, symmetric components refer to the sign–flipped versions of a frequency vector, i.e., $(\pm n_{1},\pm n_{2},\dots,\pm n_{m})$, which are always grouped together in the spectrum of a real–valued function. As a result, redundant terms that would otherwise appear in the complex exponential form in (5) are implicitly bundled together, allowing real–valued functions to be expressed more efficiently while still spanning the complete set of available basis functions.
@@
The classical Fourier series (5) and (6), which are defined by a rigid, evenly–spaced grid of frequencies, can be generalized into the nonharmonic Fourier series [2, 1]. This concept provides a mathematical bridge to the Fourier Transform (FT), which uses a continuous spectrum of all possible frequencies. The nonharmonic series occupies a middle ground: like a classical series, it is a discrete sum of sinusoids, but it inherits the flexibility of the FT by allowing its frequencies to be arbitrary real numbers. This transformation is formally achieved by replacing the fixed set of integer frequency vectors, $\mathbb{N}_{0}^{m}$, with a general finite (for practical implementation) set of real–valued vectors, $\mathcal{N}=\{\boldsymbol{n}^{1},\boldsymbol{n}^{2},\dots,\boldsymbol{n}^{N}\}$. This flexibility allows for the construction of an adaptive, task-specific Fourier basis capable of efficiently capturing the dominant spectral features of both periodic and nonperiodic functions. From this point forward, unless specified otherwise, we will use the term “Fourier series” to refer to this more general nonharmonic form, where the frequencies are from the set $\mathcal{N}$. The classical harmonic series will be considered a special case for when $\mathcal{N}\subseteq\mathbb{N}_{0}^{m}$.
@@
With this change of frequency set just described, the classical series defined in Equations (5) and (6) are generalized to their nonharmonic forms. The complex exponential form (5) is generalized using a frequency set $\mathcal{N}\subseteq\mathbb{R}^{m}$ as
$\widehat{f}_{NH}(\boldsymbol{x})=\sum_{\boldsymbol{n}\in\mathcal{N}}c_{\boldsymbol{n}}e^{i\boldsymbol{n}\cdot\boldsymbol{x}},$ (7)
@@
and the real trigonometric separable form (6) is generalized using a frequency set of nonnegative real vectors, $\mathcal{N}\subseteq\mathbb{R}_{\geq 0}^{m}$ as
$\widehat{f}_{NH}(\boldsymbol{x})=\sum_{\boldsymbol{n}\in\mathcal{N}}\Bigl{[}$ $a_{1}^{(\boldsymbol{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\cos(n_{m}x_{m})$ (8)
$+$ $a_{2}^{(\boldsymbol{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\sin(n_{m}x_{m})+\ \dots$
$+$ $a_{2^{m}}^{(\boldsymbol{n})}\,\sin(n_{1}x_{1})\sin(n_{2}x_{2})\dots\sin(n_{m}x_{m})\Bigr{]}.$
@@
While both (7) and (8) represent standard forms of the Fourier series, they are not readily compatible with NN architectures: in the exponential form (7), the basis functions are complex exponentials, which require complex–valued weights, activations, and gradient updates. Since traditional NNs are designed for real–valued computation, this would necessitate specialized handling or an explicit separation into real and imaginary parts, increasing implementation complexity. Additionally, the trigonometric form (8) expresses each multidimensional basis as a product of one–dimensional sinesine and cosinecosine factors. Thus, these bases become difficult to be represented by traditional NNs, which combine inputs additively through weighted sums before applying nonlinear activations.
To resolve the representation issues above, we opt for the cosinecosine phase–shifted form, where, by phase shift we mean the translation of a cosinecosine wave along its direction of oscillation. By converting the Fourier basis into a sum of cosinescosines with phase shifts, we eliminate the need for complex numbers in (7) and avoid the multiplicative interactions present in the trigonometric form of (8). This makes the resulting representation more compatible with the additive structure of traditional, feedforward NNs. Next, we rewrite the standard separable trigonometric Fourier series in a cosinecosine phase–shifted form, beginning with the two–dimensional case before generalizing to $m$ dimensions. We also show the relationship between the Fourier coefficients in the separable trigonometric form and the amplitudes and phase shifts in the cosinecosine phase–shifted form.
@@ 3.1 Cosine Phase–Shifted Basis for Two–dimensional Input
$\widehat{f}_{NH}(x_{1},x_{2})=\sum_{\boldsymbol{n}\in\mathcal{N}}\Big{[}$ $a_{1}^{(\boldsymbol{n})}\cos(n_{1}x_{1})\cos(n_{2}x_{2})+a_{2}^{(\boldsymbol{n})}\cos(n_{1}x_{1})\sin(n_{2}x_{2})$ (9)
$+$ $\ a_{3}^{(\boldsymbol{n})}\sin(n_{1}x_{1})\cos(n_{2}x_{2})+a_{4}^{(\boldsymbol{n})}\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Big{]},$
@@
where $\boldsymbol{n}=(n_{1},n_{2})$. Alternatively, (9) can be expressed in the cosinecosine phase-shift form
$\widehat{f}_{NH}(x_{1},x_{2})$ $=\sum_{\boldsymbol{n}\in\mathcal{N}}\left[A^{(\boldsymbol{n})}_{1}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\boldsymbol{n})}\right)+A^{(\boldsymbol{n})}_{2}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\boldsymbol{n})}\right)\right]$ (10)
$=\sum_{\boldsymbol{n}\in\mathcal{N}}H_{\boldsymbol{n}},$
@@
where $H_{\boldsymbol{n}}=A^{(\boldsymbol{n})}_{1}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\boldsymbol{n})}\right)+A^{(\boldsymbol{n})}_{2}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\boldsymbol{n})}\right)$ can be viewed as an implicit linear combination of the four separable Fourier basis functions associated with $\boldsymbol{n}$. $A^{(\boldsymbol{n})}_{1}$ and $A^{(\boldsymbol{n})}_{2}$ are the amplitudes of the phase–shifted cosinescosines, and $\phi_{1}$, $\phi_{2}$ are the respective phase shifts. The equivalence of (9) and (10) is shown in Appendix A with the detailed derivation of the relation that expresses the Fourier coefficients $a_{1}^{(\boldsymbol{n})}$, $a_{2}^{(\boldsymbol{n})}$, $a_{3}^{(\boldsymbol{n})}$ and $a_{4}^{(\boldsymbol{n})}$ in terms of the amplitudes and phase shifts as
$a_{1}^{(\boldsymbol{n})}$ $=A^{(\boldsymbol{n})}_{1}\cos\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\cos\phi_{2}^{(\boldsymbol{n})},\quad a_{2}^{(\boldsymbol{n})}=A^{\boldsymbol{(n)}}_{1}\sin\phi_{1}^{(\boldsymbol{n})}-A^{(\boldsymbol{n})}_{2}\sin\phi_{2}^{(\boldsymbol{n})},$ (11)
$a_{3}^{(\boldsymbol{n})}$ $=A^{(\boldsymbol{n})}_{1}\sin\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\sin\phi_{2}^{(\boldsymbol{n})},\quad a_{4}^{(\boldsymbol{n})}=-A^{(\boldsymbol{n})}_{1}\cos\phi_{1}^{(\boldsymbol{n})}+A^{(\boldsymbol{n})}_{2}\cos\phi_{2}^{(\boldsymbol{n})}.$
@@
Notice how the linear combination of two phase–shifted cosinecosine bases in (10) is implicitly generating the linear combination of the four separable trigonometric bases in (9), both associated with same frequency components. The same frequency pair $(n_{1},n_{2})$ occurs inside both cosinecosine bases with different amplitudes and phase–shifts, and with the sign of the second element flipped. To extend this idea of expressing separable trigonometric bases with phase–shifted cosinescosines to $m$ dimensions, we define the following matrix.
Definition 1.
@@
Let $m\geq 2$ be a positive integer and $l=2^{m-1}$. The mm–Lexi Sign Matrix $S$ is the $l\times m$ matrix whose rows are denoted by
$\boldsymbol{e}^{(i)}=\big{(}e_{1}^{(i)},e_{2}^{(i)},\dots,e_{m}^{(i)}\big{)},\quad i=1,2,\dots,l,$
@@
with $e_{1}^{(i)}=1$ and $\big{(}e_{2}^{(i)},e_{3}^{(i)},\dots,e_{m}^{(i)}\big{)}\in\{1,-1\}^{m-1}$. The rows $\boldsymbol{e}^{(i)}$ are arranged in lexicographic order with $1$ taken to have higher priority than 1-1. Explicitly,
$S=\begin{bmatrix}\boldsymbol{e}^{(1)}\\[0.0pt] \boldsymbol{e}^{(2)}\\[0.0pt] \boldsymbol{e}^{(3)}\\[0.0pt] \boldsymbol{e}^{(4)}\\[0.0pt] \vdots\\[1.0pt] \boldsymbol{e}^{(l-1)}\\[0.0pt] \boldsymbol{e}^{(l)}\end{bmatrix}=\begin{bmatrix}1&1&1&\dots&1&1\\ 1&1&1&\dots&1&-1\\ 1&1&1&\dots&-1&1\\ 1&1&1&\dots&-1&-1\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&-1&-1&\dots&-1&1\\ 1&-1&-1&\dots&-1&-1\end{bmatrix}_{l\times m}.$ (12)
For instance, if $m=3$, then
$S=\begin{bmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{bmatrix}$.
@@
$S$ helps us to construct the cosinecosine phase–shifted form in $m$-dimensions, which we discuss next.

### v3

@@ 3 Mathematical Preliminaries
$\widehat{f}_{H}(\bm{x})=\sum_{\bm{n}\in\mathbb{Z}^{m}}c_{\bm{n}}\,e^{i\bm{n}\cdot\bm{x}},$ (5)
where $\bm{x}=(x_{1},x_{2},\dots,x_{m})\in[-\pi,\pi]^{m}$ , $\bm{n}=(n_{1},n_{2},\dots,n_{m})\in\mathbb{Z}^{m}$ and $\bm{n}\cdot\bm{x}$ represents their dot product. The analytical expression for calculating the Fourier coefficients is
$c_{\bm{n}}=\frac{1}{(2\pi)^{m}}\int_{[-\pi,\pi]^{m}}f(\bm{x})\,e^{-i\bm{n}\cdot\bm{x}}\,d\bm{x},$
@@
$\widehat{f}_{H}(\bm{x})=\sum_{\bm{n}\in\mathbb{N}_{0}^{m}}\Bigl[$ $a_{1}^{(\bm{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\cos(n_{m}x_{m})$ (6)
$+$ $a_{2}^{(\bm{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\sin(n_{m}x_{m})+\ \dots$
$+$ $a_{2^{m}}^{(\bm{n})}\,\sin(n_{1}x_{1})\sin(n_{2}x_{2})\dots\sin(n_{m}x_{m})\Bigr],$
@@
for input $\bm{x}=(x_{1},x_{2},\dots,x_{m})\in[-\pi,\pi]^{m}$ and frequency components $\bm{n}=(n_{1},n_{2},\dots,n_{m})\in\mathbb{N}_{0}=\{0,1,2,\dots\}$. This trigonometric separable form has the advantage that the set of basis functions, associated with $\bm{n}$, simultaneously accounts for all possible $2^{m}$ symmetric frequency components. Here, symmetric components refer to the sign–flipped versions of a frequency vector, i.e., $(\pm n_{1},\pm n_{2},\dots,\pm n_{m})$, which are always grouped together in the spectrum of a real–valued function. As a result, redundant terms that would otherwise appear in the complex exponential form in Equation 5 are implicitly bundled together, allowing real–valued functions to be expressed more efficiently while still spanning the complete set of available basis functions.
@@
The classical Fourier series Equation 5 and Equation 6, which are defined by a rigid, evenly–spaced grid of frequencies, can be generalized into the nonharmonic Fourier series (Duffin and Schaeffer, 1952; Young, 1981). This concept provides a mathematical bridge to the Fourier Transform (FT), which uses a continuous spectrum of all possible frequencies. The nonharmonic series occupies a middle ground: like a classical series, it is a discrete sum of sinusoids, but it inherits the flexibility of the FT by allowing its frequencies to be arbitrary real numbers. This transformation is formally achieved by replacing the fixed set of integer frequency vectors, $\mathbb{N}_{0}^{m}$, with a general finite (for practical implementation) set of real–valued vectors, $\mathcal{N}=\{\bm{n}^{1},\bm{n}^{2},\dots,\bm{n}^{N}\}$. This flexibility allows for the construction of an adaptive, task-specific Fourier basis capable of efficiently capturing the dominant spectral features of both periodic and nonperiodic functions. From this point forward, unless specified otherwise, we will use the term “Fourier series” to refer to this more general nonharmonic form, where the frequencies are from the set $\mathcal{N}$. The classical harmonic series will be considered a special case for when $\mathcal{N}\subseteq\mathbb{N}_{0}^{m}$.
@@
With this change of frequency set just described, the classical series defined in Equation 5 and Equation 6 are generalized to their nonharmonic forms. The complex exponential form Equation 5 is generalized using a frequency set $\mathcal{N}\subseteq\mathbb{R}^{m}$ as
$\widehat{f}_{NH}(\bm{x})=\sum_{\bm{n}\in\mathcal{N}}c_{\bm{n}}e^{i\bm{n}\cdot\bm{x}},$ (7)
@@
and the real trigonometric separable form Equation 6 is generalized using a frequency set of nonnegative real vectors, $\mathcal{N}\subseteq\mathbb{R}_{\geq 0}^{m}$ as
$\widehat{f}_{NH}(\bm{x})=\sum_{\bm{n}\in\mathcal{N}}\Bigl[$ $a_{1}^{(\bm{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\cos(n_{m}x_{m})$ (8)
$+$ $a_{2}^{(\bm{n})}\,\cos(n_{1}x_{1})\cos(n_{2}x_{2})\dots\sin(n_{m}x_{m})+\ \dots$
$+$ $a_{2^{m}}^{(\bm{n})}\,\sin(n_{1}x_{1})\sin(n_{2}x_{2})\dots\sin(n_{m}x_{m})\Bigr].$
@@
While both Equation 7 and Equation 8 represent standard forms of the Fourier series, they are not readily compatible with NN architectures: in the exponential form Equation 7, the basis functions are complex exponentials, which require complex–valued weights, activations, and gradient updates. Since traditional NNs are designed for real–valued computation, this would necessitate specialized handling or an explicit separation into real and imaginary parts, increasing implementation complexity. Additionally, the trigonometric form Equation 8 expresses each multidimensional basis as a product of one–dimensional sinesine and cosinecosine factors. Thus, these bases become difficult to be represented by traditional NNs, which combine inputs additively through weighted sums before applying nonlinear activations.
To resolve the representation issues above, we opt for the cosinecosine phase–shifted form, where, by phase shift we mean the translation of a cosinecosine wave along its direction of oscillation. By converting the Fourier basis into a sum of cosinescosines with phase shifts, we eliminate the need for complex numbers in Equation 7 and avoid the multiplicative interactions present in the trigonometric form of Equation 8. This makes the resulting representation more compatible with the additive structure of traditional, feedforward NNs. Next, we rewrite the standard separable trigonometric Fourier series in a cosinecosine phase–shifted form, beginning with the two–dimensional case before generalizing to $m$ dimensions. We also show the relationship between the Fourier coefficients in the separable trigonometric form and the amplitudes and phase shifts in the cosinecosine phase–shifted form.
@@ 3.1 Cosine Phase–Shifted Basis for Two–dimensional Input
$\widehat{f}_{NH}(x_{1},x_{2})=\sum_{\bm{n}\in\mathcal{N}}\Big[$ $a_{1}^{(\bm{n})}\cos(n_{1}x_{1})\cos(n_{2}x_{2})+a_{2}^{(\bm{n})}\cos(n_{1}x_{1})\sin(n_{2}x_{2})$ (9)
$+$ $\ a_{3}^{(\bm{n})}\sin(n_{1}x_{1})\cos(n_{2}x_{2})+a_{4}^{(\bm{n})}\sin(n_{1}x_{1})\sin(n_{2}x_{2})\Big],$
@@
where $\bm{n}=(n_{1},n_{2})$. Alternatively, Equation 9 can be expressed in the cosinecosine phase-shift form
$\widehat{f}_{NH}(x_{1},x_{2})$ $=\sum_{\bm{n}\in\mathcal{N}}\left[A^{(\bm{n})}_{1}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\bm{n})}\right)+A^{(\bm{n})}_{2}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\bm{n})}\right)\right]$ (10)
$=\sum_{\bm{n}\in\mathcal{N}}H_{\bm{n}},$
@@
where $H_{\bm{n}}=A^{(\bm{n})}_{1}\cos\left(n_{1}x_{1}+n_{2}x_{2}-\phi_{1}^{(\bm{n})}\right)+A^{(\bm{n})}_{2}\cos\left(n_{1}x_{1}-n_{2}x_{2}-\phi_{2}^{(\bm{n})}\right)$ can be viewed as an implicit linear combination of the four separable Fourier basis functions associated with $\bm{n}$. $A^{(\bm{n})}_{1}$ and $A^{(\bm{n})}_{2}$ are the amplitudes of the phase–shifted cosinescosines, and $\phi_{1}$, $\phi_{2}$ are the respective phase shifts. The equivalence of Equation 9 and Equation 10 is shown in Appendix A with the detailed derivation of the relation that expresses the Fourier coefficients $a_{1}^{(\bm{n})}$, $a_{2}^{(\bm{n})}$, $a_{3}^{(\bm{n})}$ and $a_{4}^{(\bm{n})}$ in terms of the amplitudes and phase shifts as
$a_{1}^{(\bm{n})}$ $=A^{(\bm{n})}_{1}\cos\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\cos\phi_{2}^{(\bm{n})},\quad a_{2}^{(\bm{n})}=A^{\bm{(n)}}_{1}\sin\phi_{1}^{(\bm{n})}-A^{(\bm{n})}_{2}\sin\phi_{2}^{(\bm{n})},$ (11)
$a_{3}^{(\bm{n})}$ $=A^{(\bm{n})}_{1}\sin\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\sin\phi_{2}^{(\bm{n})},\quad a_{4}^{(\bm{n})}=-A^{(\bm{n})}_{1}\cos\phi_{1}^{(\bm{n})}+A^{(\bm{n})}_{2}\cos\phi_{2}^{(\bm{n})}.$
@@
Notice how the linear combination of two phase–shifted cosinecosine bases in Equation 10 is implicitly generating the linear combination of the four separable trigonometric bases in Equation 9, both associated with same frequency components. The same frequency pair $(n_{1},n_{2})$ occurs inside both cosinecosine bases with different amplitudes and phase–shifts, and with the sign of the second element flipped. To extend this idea of expressing separable trigonometric bases with phase–shifted cosinescosines to $m$ dimensions, we define the following matrix.
Definition 1 (mm–Lexi Sign Matrix)
@@
Let $m\geq 2$ be a positive integer and $l=2^{m-1}$. The mm–Lexi Sign Matrix $S^{(m)}$ is the $l\times m$ matrix whose rows are denoted by
$\bm{e}^{(i)}=\big(e_{1}^{(i)},e_{2}^{(i)},\dots,e_{m}^{(i)}\big),\quad i=1,2,\dots,l,$
@@
with $e_{1}^{(i)}=1$ and $\big(e_{2}^{(i)},e_{3}^{(i)},\dots,e_{m}^{(i)}\big)\in\{1,-1\}^{m-1}$, i.e., the (m1)(m-1)–fold Cartesian product of the set $\{1,-1\}$. The rows $\bm{e}^{(i)}$ are arranged in lexicographic order with $1$ taken to have higher priority than 1-1. Explicitly,
$S^{(m)}=\begin{bmatrix}\bm{e}^{(1)}\\[0.0pt] \bm{e}^{(2)}\\[0.0pt] \bm{e}^{(3)}\\[0.0pt] \bm{e}^{(4)}\\[0.0pt] \vdots\\[1.0pt] \bm{e}^{(l-1)}\\[0.0pt] \bm{e}^{(l)}\end{bmatrix}=\begin{bmatrix}1&1&1&\dots&1&1\\ 1&1&1&\dots&1&-1\\ 1&1&1&\dots&-1&1\\ 1&1&1&\dots&-1&-1\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&-1&-1&\dots&-1&1\\ 1&-1&-1&\dots&-1&-1\end{bmatrix}_{l\times m}.$ (12)
For instance, for the cases of $m=2$ and $m=3$, we have
$S^{(2)}=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix},\quad S^{(3)}=\begin{bmatrix}1&1&1\\ 1&1&-1\\ 1&-1&1\\ 1&-1&-1\end{bmatrix}$.
@@
$S^{(m)}$ helps us to construct the separable bases with phase–shifted cosinescosines in $m$-dimensions, which we discuss next.